### v2

# ARES: Alternating Reinforcement Learning and Supervised Fine-Tuning for Enhanced Multi-Modal Chain-of-Thought Reasoning Through Diverse AI Feedback

Abstract
Large Multimodal Models (LMMs) excel at comprehending human instructions and demonstrate remarkable results across a broad spectrum of tasks. Reinforcement Learning from Human Feedback (RLHF) and AI Feedback (RLAIF) further refine LLMs by aligning them with specific preferences. These methods primarily use ranking-based feedback for entire generations. With advanced AI models (Teacher), such as GPT-4 and Claude 3 Opus, we can request various types of detailed feedback that are expensive for humans to provide. We propose a two-stage algorithm ARES that Alternates REinforcement Learning (RL) and Supervised Fine-Tuning (SFT). First, we ask the Teacher to score how much each sentence contributes to solving the problem in a Chain-of-Thought (CoT). This sentence-level feedback allows us to consider individual valuable segments, providing more granular rewards for the RL procedure. Second, we ask the Teacher to correct wrong reasoning after the RL stage. The RL procedure requires substantial hyperparameter tuning and often generates errors such as repetitive words and incomplete sentences. With correction feedback, we stabilize the RL fine-tuned model through SFT. We conduct experiments on the multi-modal datasets ScienceQA and A-OKVQA to demonstrate the effectiveness of our proposal. The ARES rationale achieves around 70% win rate compared to baseline models judged by GPT-4o. Additionally, we observe that the improved rationale reasoning leads to a 2.5% increase in inference answer accuracy on average for the multi-modal datasets. 111Code: https://github.com/Amyyyyeah/ARES
ARES: Alternating Reinforcement Learning and Supervised Fine-Tuning for Enhanced Multi-Modal Chain-of-Thought Reasoning
Through Diverse AI Feedback
1 Introduction
Large Language Models (LLMs) and Large Multimodal Models demonstrate remarkable performance across diverse language and multi-modal tasks (Brown et al., 2020; Chowdhery et al., 2022; Touvron et al., 2023; Zhang et al., 2022a; Liu et al., 2023a; Goyal et al., 2023). However, these Large Models (LMs) often generate toxic and biased content (Gehman et al., 2020; Tamkin et al., 2021) because LMs are primarily trained to predict the next token based on extensive corpus datasets. To align LM behavior more closely with user preferences, previous works (Glaese et al., 2022; Ouyang et al., 2022) fine-tune their models using Reinforcement Learning from Human Feedback (RLHF). Furthermore, with advancements in LMs, advanced LM feedback can replace costly human feedback, yielding Reinforcement Learning from AI Feedback (RLAIF) (Lee et al., 2023; Bai et al., 2022; Yuan et al., 2024).
However, RLHF and RLAIF encounter two significant challenges. First, both methods often utilize ranking-based feedback (Ouyang et al., 2022), which orders the generated samples by preference. For instance, if sample is preferred over sample , the model is fine-tuned to generate more outputs like and fewer like . However, if contains certain valuable parts, they may be discarded. To alleviate these issues, Lightman et al. (2023); Luo et al. (2024) propose sentence-level feedback, applying it solely to the reward model without Reinforcement Learning (RL), called the Process-supervised Reward Model (PRM). It demonstrates its potential through search algorithms on the PRM, such as best-of- or Monte Carlo Tree Search (MCTS). Furthermore, Wang et al. (2024) demonstrate the effectiveness of RL with sentence-level feedback by heuristically scoring each sentence in math problems, where evaluating the predicted answer is straightforward. Thus, sentence-level feedback exhibits significant promise compared to existing ranking-based feedback. Nonetheless, acquiring sentence-level feedback is more costly than ranking-based feedback.
Second, the RL process is inherently unstable and requires extensive hyperparameter tuning (Eimer et al., 2023). This instability often results in the generation of repetitive words and truncated sentences. Hyperparameter tuning becomes an enormous burden as the model size increases, making exhaustive tuning seemingly impossible, especially for individuals. The existing RLHF method (Ouyang et al., 2022) recycles the dataset used in pretraining within the loss function with Proximal Policy Optimization (PPO) (Schulman et al., 2017) to mitigate this degeneration problem. However, this approach prevents the model from fully maximizing the sum of rewards through RL and may limit the opportunity for diverse improvements, which differ from the pretraining distribution.
In this work, we aim to address the two challenges mentioned above through various types of feedback using an advanced AI model as a Teacher. Many advanced AI models, including GPT-4 and Claude 3 Opus, are already used as evaluators for many tasks and generate reliable human-level answers (Liu et al., 2023b; Sottana et al., 2023). We request a score from the Teacher for each sentence, ranging from to . Each score indicates how much a sentence contributes to solving the problem. This provides detailed reward feedback to the training model and can be applied to both mathematical and more general multi-modal Chain-of-Thought (CoT) (or rationale reasoning) problems. We ask the Teacher to identify and correct minor errors in the RL results, such as incorrect or cut-off parts. With this corrected dataset, we fine-tune the model using Supervised Fine-Tuning (SFT). This stage allows the model to maximize the rewards while properly deviating from the pretraining distribution. In summary, we propose a hybrid algorithm ARES that Alternates REinforcement Learning and Supervised Fine-Tuning.
To evaluate how much rationale reasoning can be improved through the ARES framework, we use the ScienceQA and A-OKVQA datasets, which are large-scale, multi-modal datasets that include rationale reasoning data. We use Multimodal-CoT (MM-CoT) (Zhang et al., 2023b) as the baseline. MM-CoT employs two separate models: one model is responsible for generating rationale, and the other model, an inference model, processes the concatenated input (problem and generated rationale). This distinct framework enhances performance, even with relatively smaller models like (Chia et al., 2023) (251M) and (790M) with ViT feature (Dosovitskiy et al., 2021). We perform ARES on the rationale reasoning model of MM-CoT. We compare ARES rationale reasoning with that of MM-CoT through GPT-4o, determining which rationale is better and computing the win rate. Additionally, we check whether the improved rationale reasoning leads to better answer accuracy. Our results show that our rationale reasoning outperforms the baselines with around 70% win rate and demonstrates 2.5% increase in inference answer accuracy on average for the different model sizes and the multi-modal tasks.
In summary, our key contributions are:
We propose ARES, leveraging diverse types of feedback from advanced AI (Teacher) to enhance CoT reasoning for multi-modal tasks.
ARES properly reflects the direction in which the model is fine-tuned through RL and stabilizes the RL fine-tuning procedure with SFT.
ARES generates better rationale chains compared to baselines judged by
GPT-4oand improves inference answer accuracy for multi-modal tasks.
2 Methodology
This section briefly introduces the preliminaries in Section 2.1 and presents our two-stage hybrid algorithm ARES that Alternates REinforcement Learning and Supervised Fine-Tuning. We request a score for each sentence in the Chain-of-Thought (CoT) from the advanced AI model (Teacher) to determine how much it contributes to solving the problem (Section 2.2). We perform Reinforcement Learning (RL) with the score feedback on our training model. The Teacher corrects minor errors such as truncated or slightly incorrect sentences, thereby performing Supervised Fine-Tuning (SFT) (Section 2.2).
2.1 Preliminaries
For an input , a Transformer-based (Vaswani et al., 2023) model parameterized by generates the output composed of sentences .
where indicates previous tokens. To proceed with RL finetuning, Ouyang et al. (2022) train an outcome-supervised reward model (ORM) using ranking-based feedback. With more fine-grained feedback like sentence-level, Lightman et al. (2023); Wang et al. (2024) train a process-supervised reward model (PRM). Instead of training a reward model, we request score feedback for each sentence from an advanced AI such as GPT-4 where represents previous sentences.
2.2 Reinforcement Learning
Reinforcement Learning (RL) fine-tunes our model to maximize sum of sentence rewards from an advanced AI model such as GPT-4 and Claude 3 Opus. The RL objective is as follows:
where is a discount factor. We use Proximal Policy Optimization (PPO) (Schulman et al., 2017) to achieve this RL objective, treating sentences as actions (Equation 3).
where is the original policy (baseline model) and is an advantage estimator at timestep . PPO is commonly leveraged in Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022) and AI Feedback (RLAIF) (Bai et al., 2022). PPO’s conservative update prevents the training model from deviating too far from the original model, thus avoiding degeneration.
Sentence-Level Nuanced Feedback: We request a score between and for each sentence in CoT through the advanced AI for RL. The closer the score is to , the more relevant and helpful it is to solving the problem. Table 5 presents the prompt format. We additionally shift the reward distribution by to center it at (Zheng et al., 2023). Therefore, the actual range is from to . Using these nuanced scores, the RL fine-tuned model exhibits emergent behaviors (please refer to Section 4). This allows us to understand the direction in which the model is intended to change through RL.
Advantages of Using Advanced AI for Score Feedback: Although calling the API has disadvantages, such as incurring costs or facing usage limits, there exist several advantages to using the advanced AI for feedback. First, there is no need to train a reward model. Second, as the RL fine-tuned model begins to generate out-of-distribution outputs that differ from the data used to train the reward model, it becomes challenging for the trained reward model to provide accurate rewards. However, this out-of-distribution problem is effectively addressed with the advanced AI.
RL Challenge: One of the challenging factors for RL is hyperparameter tuning (Eimer et al., 2023). This often results in generating repetitive words and truncated sentences (Ouyang et al., 2022). Additionally, as the model size increases, finding working hyperparameters becomes infeasible for individuals. To alleviate this issue, we utilize correction feedback from the advanced AI as the second stage (Section 2.3), and proceed with the supervised fine-tuning to stabilize the RL fine-tuned model.
2.3 Correction: Supervised Fine-Tuning
The RL fine-tuning procedure makes model changes to maximize the reward sum, such as correcting mistakes or explaining why other options cannot be the answer. However, without highly tuned hyperparameters (Eimer et al., 2023), the model after the RL phase may result in errors such as repeated sentences, truncated sentences, or incorrect content for some data points. (See examples in Appendix D.)
Correction Feedback: Given the success of LLMs and LMMs in a wide range of areas (Brown et al., 2020; Chowdhery et al., 2022; Zhang et al., 2022a), we are not restricted to requesting feedback in the form of scores. We request correction feedback from advanced AI (Teacher) for sentences containing errors after the RL process, and obtain a corrected dataset . Since the supervised fine-tuning is more stable and finding appropriate hyperparameters is easier than RL, we proceed with supervised fine-tuning using exactly as in common autoregressive model (Vaswani et al., 2023) training to stabilize the RL fine-tuned model. This reduces the burden of RL’s exhaustive hyperparameter tuning and properly guides the direction in which the training model wants to change.
How Correction Feedback Helps RL: RL increases the probability of positively rewarded actions (or sentences) and decreases the probability for negative rewards. The direction of learning is determined by the reward (scalar) value. However, the opposite direction of the reward is sometimes required. For example, suppose there is a truncated sentence in CoT. gets a negative score because it is an incomplete sentence (Table 13). If there is no correction stage, the probability of is simply reduced. What if contains some valuable part? This valuable part is ignored, and its probability decreases. To alleviate this issue, we instead receive the corrected sentence as feedback and encourage the training model to generate complete sentences, which is very challenging to achieve with only RL. Table 16 shows more examples of how the correction stage helps the RL stage by maintaining the reasoning context while changing the erroneous parts.
Additionally, RL is primarily fine-tuned through PPO (Schulman et al., 2017) to prevent the model from deviating too much from the original model. The KL divergence penalty further prevents deviation. However, this penalty often causes the model’s degeneration. As a solution, InstructGPT (Ouyang et al., 2022) proposes PPO-ptx, where the supervised fine-tuning term with the pretraining dataset is included in the loss function. While this aims to align the training model with specific preferences, it tends to anchor the model to the pretraining dataset. Instead, we perform supervised fine-tuning through the Teacher’s correction feedback to allow the training model to more freely adapt and meet specific preferences without degeneration.
2.4 Algorithm Details
We propose a hybrid algorithm Alternating between REinforcement learning and Supervised fine-tuning (ARES). Figure 2 illustrates the ARES pipeline. First, we prepare a model with a given training dataset and generate rationale reasoning composed of several sentences for input. For the RL procedure to align the training model with a preference, we request scores for each sentence. The RL result may include some incorrect parts (colored in red as the 4th sentence in the RL result box), but it aims to maximize the rewards provided. Next, we request correction feedback and create a corrected dataset (colored in blue as the 3rd sentence in the Corrected Rationale box) for the supervised fine-tuning stage. We then repeat the process from the RL stage until convergence.
3 Experimental Setup
Data: We first evaluate our proposed method on the ScienceQA (Lu et al., 2022a) dataset, a large-scale, multi-modal science dataset designed to assess multi-hop reasoning abilities. We choose ScienceQA because it contains reasoning chains to derive the answer. Each problem consists of a question, multiple options, multi-modal contexts, a correct answer, and an annotated lecture or solution chain (note that around lack the solution chain). In addition, we conduct experiments on A-OKVQA (Schwenk et al., 2022), a knowledge-based multi-modal benchmark with a diverse set of challenging questions paired with rationales, demanding non-trivial commonsense knowledge (see Appendix B).
Baselines: We mainly compare our method with Multimodal-CoT (MM-CoT) (Zhang et al., 2023b) as the baseline because it utilizes reasoning chains to solve multi-modal tasks. MM-CoT leverages two distinct models: the first generates a rationale for a given problem, and the second, an inference model, takes the concatenated input (problem and generated rationale). This separated framework shows improved performance, even for relatively small models such as (Chia et al., 2023) (M) and (M). We use the rationale model provided by MM-CoT for ScienceQA and retrain the rationale model ourselves for A-OKVQA because there is no provided model.
Prompts for Feedback: Since our proposed ARES requests different types of feedback for each stage, a corresponding prompt exists separately. We use Claude 3 Haiku for all training to get feedback because it is approximately times cheaper than the top competing models, yet still demonstrates decent performance. We first request scores ranging from to for each sentence in CoT to proceed with the RL stage. To obtain reasonable scores, we let Haiku consider the starting point of thought, the process of elimination, or true statements. (See Table 5.)
In order to collect the corrected dataset for the SFT stage, we let Haiku refer to the given problem and correct the answer as the prompt. We ask Haiku to maintain the format of the existing rationale chains as much as possible and correct only the parts that require correction. The RL stage often makes the training model generate repetitive sentences. This repetition is not easily removed even by GPT-4 when the repetitive sentence exists in the middle of rationale reasoning. To reduce the burden of feedback, we simply hard-code the removal of repetitive sentences before adding the generated rationale to the prompt. (See Appendix C.2.)
Training Details: For the RL stage, we use a learning rate of and epochs for PPO with a batch size of for both ScienceQA and A-OKVQA. The learning rate for is with epochs for PPO and a batch size of for both tasks. We proceed with 2 rounds of our pipeline for and 2 rounds for for ScienceQA. For A-OKVQA, we proceed with 1 round for both model sizes. For the SFT stage for correction, we follow the hyperparameters used in MM-CoT for both model sizes. Additionally, we replace MM-CoT’s inference model, which is the same size as the rationale model, with the Low-Rank Adaptation (LoRA) (Hu et al., 2021) added to the rationale model (Figure 4). The LoRA adapter effectively utilizes the rationale model’s features with a small number of weights, enabling 2x–14x faster inference compared to MM-CoT, which introduces a separate inference model (See the time comparison in Table 7 and Table 8). For more detailed settings, please refer to Appendix C.
Evaluation Metrics: We use two main metrics to test how our pipeline (ARES) improves rationale reasoning quality. First, we evaluate ARES’s rationale reasoning quality against baseline models since we enhance our model based on them. For two different model sizes ( and ) and two tasks (ScienceQA and A-OKVQA), rationale reasoning quality is evaluated by GPT-4o-2024-05-13 and the win rate is calculated (Section 4.3). The GPT-4 series is actively used as an evaluation metric, replacing human judgment for various domains (Liu et al., 2023b; Sottana et al., 2023). Second, we assess how the improved rationale reasoning impacts answer accuracy (Section 4.4). This evaluation is also performed on both model sizes and tasks. Additionally, we analyze how the RL stage fine-tunes the training model and maximizes the sum of rewards in Section 4.1.
4 Experimental Results
In this section, we evaluate our pipeline ARES that Alternates REinforcement Learning and Supervised Fine-Tuning by requesting diverse types of feedback for an advanced AI model (Teacher) (Claude 3 Haiku). The goal of ARES is to improve the rationale reasoning quality. We demonstrate how ARES enhances rationale reasoning in the following sections.
4.1 Emergent Behavior Through RL
Through RL, a training model is aligned to a specific preference. Essentially, the model increases the probability of helpful sentences receiving good rewards and reduces the probability of incorrect or meaningless sentences. However, this process produces some interesting additional results.
First, it supplements rationale reasoning for some problems where rationale reasoning is insufficient. In particular, 9.5% of problems in ScienceQA have empty rationale reasoning (solution) data. The model generates nothing before the RL stage for these problems but starts generating reasoning chains afterward (See Table 14). We observe this especially when utilizing PPO’s advantage normalization or when the learning rate is large.
Second, the training model begins to explain why other options are not the answer (See Table 15). The process of elimination is a useful method for deriving answers when options are given.
4.2 Guide RL with Correction
Despite the benefits of RL, hyperparameter tuning often requires massive effort. Without meticulous tuning, the RL fine-tuned model may produce errors such as repetitive or incomplete sentences. To address these issues, we add a supervised fine-tuning (SFT) stage after RL to correct these errors. SFT is more stable than RL. We evaluate how well the SFT stage corrects errors caused by the RL stage for various RL hyperparameters. We test various RL hyperparameters such as learning rate = {5e-6, 1e-5, 2e-5, 5e-5}, batch size = {2, 4, 8, 16, 32}, and PPO epoch = {5, 10, 15}. As a result of RL, we observe that some of the sentences in rationale chains are repetitive or truncated (see Table 13 and 12). The SFT stage, with correction feedback, reflects the direction in which the model is fine-tuned through RL and appropriately guides it (Table 13 and 16). However, excessive RL learning rates or epochs cause serious degeneration of the model, such as producing no output or generating strange words, and the results of correction feedback are also unreasonable.
ScienceQA Win Rate vs 69.76% vs 73.76% A-OKVQA Win Rate vs 69.11% vs 66.96%
Model Size NAT SOC LAN TXT IMG NO G1-6 G7-12 Avg Human - 90.23 84.97 87.48 89.60 87.50 88.10 91.59 82.42 88.40 MCAN (Yu et al., 2019) 95M 56.08 46.23 58.09 59.43 51.17 55.40 51.65 59.72 54.54 Top-Down (Anderson et al., 2018) 70M 59.50 54.33 61.82 62.90 54.88 59.79 57.27 62.16 59.02 BAN (Kim et al., 2018) 112M 60.88 46.57 66.64 62.61 52.60 65.51 56.83 63.94 59.37 DFAF (Peng et al., 2019) 74M 64.03 48.82 63.55 65.88 54.49 64.11 57.12 67.17 60.72 ViLT (Kim et al., 2021) 113M 60.48 63.89 60.27 63.20 61.38 57.00 60.72 61.90 61.14 Patch-TRM (Lu et al., 2022b) 90M 65.19 46.79 65.55 66.96 55.28 64.95 58.04 67.50 61.42 VisualBERT (Li et al., 2019) 111M 59.33 69.18 61.18 62.71 62.17 58.54 62.96 59.92 61.87 UnifiedQABase (Khashabi et al., 2020) 223M 68.16 69.18 74.91 63.78 61.38 77.84 72.98 65.00 70.12 UnifiedQABase w/ CoT (Lu et al., 2022a) 223M 71.00 76.04 78.91 66.42 66.53 81.81 77.06 68.82 74.11 LLaMA-Adapter (Zhang et al., 2023a) 6B 84.37 88.30 84.36 83.72 80.32 86.90 85.83 84.05 85.19 LLaVA (Liu et al., 2023a) 13B 90.36 95.95* 88.00 89.49 88.00 90.66 90.93 90.90* 90.92 InstructBLIP (Dai et al., 2023) 11B - - - - 90.70* - - - - (Zhang et al., 2023b) 251M+251M 84.59 92.46 83.45 83.87 83.29 85.64 86.34 85.23 85.95 (Ours) 251M+30M 87.92 92.58 85.91 86.61 85.82 88.36 88.88 87.48 88.38 (Zhang et al., 2023b) 790M+790M 90.76 93.59 86.55 89.69 87.85 89.55 90.90 89.12 90.26 (Ours) 790M+76M 91.21* 92.80 89.45* 90.27* 88.35 91.22* 91.48* 90.38 91.09*
4.3 Rationale Reasoning Comparison
We check whether ARES improves the quality of rationale reasoning compared to the baseline model. GPT-4o evaluates which rationale chain is better between the rationale generated by ARES and the rationale generated by the baseline model. We randomly shuffle the rationale chains and provide them as Option A and Option B (see Appendix A.3) for a fair evaluation (Yu et al., 2023). We conduct our experiments with two different model sizes, Flan-Base and Flan-Large with ViT feature, on ScienceQA and A-OKVQA. Table 1 shows that ARES achieves around 70% win rate against each corresponding baseline model for both datasets.
4.4 Inference Accuracy
We investigate whether the improved rationale also contributes to answer inference accuracy. Table 2 shows the main results of answer inference on the ScienceQA. We evaluate our base model against the MM-CoT baseline. achieves a 2.79% improvement compared to the corresponding baseline (). The large model () shows some minimal improvement compared to the corresponding baseline. However, it’s worth noting that despite this seemingly small gain, beats B LLaVA (Liu et al., 2023a). This minimal improvement may be due to the 9.5% of ScienceQA problems needing more rationale reasoning (around 9.5% problems have empty rationale reasoning). The RL stages can only eliminate some empty rationale reasoning, which requires numerous ARES pipeline rounds. Above all, our main goal is to assess how the RL stage works and how the SFT stage aids RL.
Table 3 shows the results of answer inference on the A-OKVQA. We retrain and and evaluate these on the validation set as in (Zhang et al., 2023b) because the test set is hidden. In our experiments, MM-CoT models perform around 10% better than the reported accuracy in (Zhang et al., 2023b). ARES achieves 4.45% gains against and 2.35% for .
In addition, we demonstrate that two stages, RL and SFT, are essential through an ablation study. Figure 3 shows the rationale reasoning for 4 cases. The baseline model (MM-CoT) produces the same rationale reasoning as the dataset. However, the corrected reasoning for MM-CoT without the RL stage has insufficient information compared to the reasoning of ARES that performs RL (refer to Table 17 for more examples). Table 4 also shows that inference accuracy gradually improves as each part of ARES is executed. 1st RL indicates a single RL run on MM-CoT, and 1st ARES means one round of the ARES pipeline. 1st ARES & 2nd RL represents the second RL on 1st ARES, and finally, 2nd ARES refers to two rounds of ARES.
Model Accuracy IPVR (OPT-66B) 48.6 ViLBERT 49.1 60.96 (Ours) 65.41 65.68 (Ours) 68.03
5 Related Work
Chain-of-Thought (CoT) is a multi-step reasoning method for problem-solving that encourages LLMs to consider the intermediate reasoning steps. Zero-Shot-CoT (Kojima et al., 2023) promotes CoT by using prompts such as "Let’s think step by step" for LLMs. For Few-Shot-CoT (Zhang et al., 2022b; Wei et al., 2023), a few examples with reasoning processes are provided, allowing the model to refer to these examples and understand how to perform CoT. Wei et al. (2023) reveal that this CoT technique positively impacts the performance of large models (B), but has minimal effect on smaller models. MM-CoT (Zhang et al., 2023b) suggest that CoT is beneficial even for relatively small models, such as 200M, if the model that generates intermediate reasoning and the model that infers the answer are separated. We find that simply adding a LoRA adapter (Hu et al., 2021) to the reasoning model results in comparable performance. This framework enables the LoRA adapter to effectively utilize all features, from raw text to latent features, and generates answers 2x–14x faster than MM-CoT, which uses a separate inference model (See Table 7 and Table 8). This speed advantage arises from the fact that our framework does not require a rationale as input, whereas the separate inference model framework must first generate the rationale before using it as input.
Reinforcement Learning from Human Feedback (RLHF) (Glaese et al., 2022; Ouyang et al., 2022) and AI Feedback (RLAIF) (Bai et al., 2022) align LLMs with user preferences. Ouyang et al. (2022) collects ranked feedback from human labelers and uses this feedback to perform Reinforcement Learning (RL). Constitutional AI (CAI) (Bai et al., 2022) collects ranked AI feedback rather than costly human feedback and handles harmfulness with RL. Both approaches learn outcome-supervised reward models (ORM) using ranking-based feedback. Lightman et al. (2023), instead, propose a process-supervised reward model (PRM) that leverages sentence-level feedback for CoT. Lightman et al. (2023); Luo et al. (2024) evaluate each trained ORM and PRM with searching algorithms such as best-of- or Monte Carlo Tree Search (MCTS) by selecting the highest-scored solution, demonstrating that the PRM-selected solution outperforms the ORM-selected one. Wang et al. (2024) perform RL using PRM, providing heuristic sentence-level scores for math problems that are simple to grade. As an LLM is trained with RL and starts generating outputs different from the original distribution, these reward models would not correctly provide rewards (Pitis, 2023; Byun and Perrault, 2024). Instead of training a reward model for a more general task, we perform RL by requesting sentence-level rewards from advanced AI models such as GPT-4.
Model Accuracy 85.95 1st RL 86.70 1st ARES 87.81 1st ARES & 2nd RL 87.88 2nd ARES 88.38
6 Conclusion
We propose a hybrid algorithm, ARES, which Alternates REinforcement Learning (RL) and Supervised Fine-Tuning (SFT) to enhance multi-modal rationale reasoning for ScienceQA and A-OKVQA. ARES leverages two types of feedback: ARES requests a score from a Teacher (we used Claude 3 Haiku) for sentence-level nuanced feedback and proceeds with RL. ARES requests the Teacher to correct rationale chains after RL, stabilizing the RL fine-tuned model with SFT. ARES is designed to aid the RL procedure without massive hyperparameter tuning while properly reflecting the desired changes from RL. We evaluate the improvement in rationale reasoning produced by ARES compared to baselines using GPT-4o, and assess how much the improved rationale chains enhance inference accuracy for the two multi-modal tasks. We hope our work inspires further research on utilizing various types of AI feedback.
Limitations
Although we address general multi-modal rationale models beyond mathematical problems, receiving feedback from AI models still needs to be more reliable for more complex tasks such as graduate-level math or expert-level knowledge. For instance, some A-OKVQA problems even contain challenging questions requiring external knowledge beyond the image alone. This challenge highlights the necessity for future research to develop methods that can effectively incorporate external knowledge sources into the model. Additionally, if the model is not publicly available for free, using the API incurs costs, and there are daily usage limits.
Acknowledgments
We sincerely appreciate the generous support of computational resources provided by the Ohio Supercomputer Center (Center, 1987).
Appendix A Prompts
| Feedback | Prompt Structure |
Sentence-Level
Nuanced Feedback |
[Prompt when Image is provided]
There exists a set comprising Image, Options, Hint, and Answer for a Question. The reasoning process used to deduce the answer is provided in JSON format. Fill in "xxx" with values ranging from 0.0 to 1.0, in increments of 0.1. The reasoning may include the starting point of thought, the process of elimination, or true statements, although these may not appear to be directly related to the answer at first glance. A value closer to 0.0 indicates a completely incorrect rationale, 0.5 indicates a neutral rationale such as the initial thought process or true statements that guide later guesses towards the answer, and a value closer to 1.0 denotes a correct or relevant rationale for the question. Please just fill the "xxx" parts and only return the JSON format. If a sentence is repetitive (appeared before), then give 0.0. Question: <Question> Options: <Choices> Hint: <Hint> Answer: <Answer> { "<Rationale 1>": xxx, "<Rationale 2>": xxx, "<Rationale 3>": xxx } [Prompt when no Image is provided] There exists a set comprising Options, Hint, and Answer for a Question. The reasoning process … <same as the prompt when the Image is provided> |
| Evaluation | Prompt Structure |
| Win Rate |
You are given two rationales options (A or B). Your job is to select the better rationale between A and B for solving the given problem with the given Image, Choices, Hint, and Answer. Please output only A or B.
Question: <Question> Choices: <Choices> Hint: <Hint> Answer: <Answer> OPTION A: <Rationales> OPTION B: <Rationales> |
A.1 Prompt for Sentence-Level Nuanced Feedback
The prompt for obtaining sentence-level nuanced feedback by Claude 3 Haiku is illustrated in Table 5. Each reasoning sentence is assigned a value between and .
Values close to indicate completely incorrect rationales.
A value of represents a neutral rationale, such as an initial thought process or true statements that aid in guiding guesses towards the correct answer.
Values close to denote a correct or highly relevant rationale.
These scores enable our model to discern the direction of changes through Reinforcement Learning (RL), reflecting the extent to which a sentence aids in resolving the problem.
A.2 Prompt for Correction Feedback
Due to the challenges mentioned in Section 2.2, we adopt the correction feedback approach. The following are the specific instructions for obtaining correction feedback using Claude 3 Haiku. We have established the following seven rules for obtaining correction feedback using Claude 3 Haiku: The prompt is presented as a Table 9.
A.3 Prompt for Win Rate Evaluation
We prompt GPT-4o (2024-05-13) to choose which generated rationale is better for solving the question because we don’t have gold rationales.
Given two generated rationales (e.g., and ), we ask GPT-4o: "You are given two rationale options (A or B). Your job is to select the better rationale between A and B for solving the given problem with the given image, choices, hint, and answer. Please output only ’A’ or ’B’." (see Table 6). Yu et al. (2023) find that ChatGPT is skewed towards choosing option A, so we randomly swap options A and B for each evaluation to avoid bias.
Appendix B Difficulties with External Knowledge in A-OKVQA
The A-OKVQA dataset includes challenging questions paired with rationales that demand knowledge beyond the information available in the image. These questions cannot be answered simply by querying a knowledge base, as they require a deeper understanding and integration of external knowledge.
Our model faces difficulties with problems that cannot be resolved using only the information from the image. While our approach is designed to improve rationales by addressing grammatical errors and incomplete or incorrect statements, it struggles with questions that necessitate external knowledge.
Table 10 illustrates an example where the question requires knowledge about the typical PSI range for bicycle tires. This information is not visually apparent from the image of the bicycle alone. To answer this question correctly, one needs external knowledge about standard bicycle maintenance practices and the recommended PSI ranges for different types of bicycle tires. This highlights a challenge for our model, as it must provide correct rationales and answers without access to such external knowledge, relying solely on the image and its internal knowledge base.
In the second example (Table 11), understanding what the first two numbers of an identification tag denote requires specific external knowledge about livestock tagging systems. Such tagging systems often use coded information, where the numbers might represent the birth month, year, or other identification details. This information is not visually apparent from the image and requires familiarity with agricultural practices or livestock management, illustrating the challenge for our model in providing correct rationales and answers without external knowledge.
Appendix C Training Details
ScienceQA has K multi-modal problems, with K for training, K for validation, and K for testing. It also includes various difficulty levels from elementary to high school, covering domains like natural science, language science, and social science. In addition, we conduct experiments on A-OKVQA (Schwenk et al., 2022), a knowledge-based multi-modal benchmark with a diverse set of 25K questions A-OKVQA includes 25K questions (K for training, K for validation, and K for testing).
We adapt the same T5 encoder-decoder architecture (Raffel et al., 2023) under Base and Large settings, following (Zhang et al., 2023b), and initialized with Flan-Alpaca (Chia et al., 2023).
C.1 Reinforcement Learning
For the and training on the ScienceQA and A-OKVQA dataset, we employ the following settings:
Common Settings: We use top- sampling with and sample 4 actions. The initial coefficient for the Kullback-Leibler (KL) divergence is set to . The range for clipping the probability ratios in PPO is . The discount factor is set to . Token length is constrained to . We train the model using 4 NVIDIA A100 80GB GPUs.
Specific Settings: We use a learning rate of and epochs for PPO with a batch size of . Advantage normalization is applied for , and gradient accumulation steps are set to .
Specific Settings: The learning rate for is with 5 epochs for PPO and a batch size of for both tasks. Advantage normalization is not used and gradient accumulation steps are set to .
C.2 Supervised Fine-Tuning
We use a batch size of 8 and train for 20 epochs with a learning rate of for , following (Zhang et al., 2023b). For , we use a batch size of and train for epochs with a learning rate of . The output length is set to tokens. Training for utilizes 1 A100 GPU, while training for utilizes 4 A100 GPUs. In the MM-CoT paper (Zhang et al., 2023b), because the final_eval setting was not consistent, we retrained the base model with final_eval=true and the large model with final_eval=false for consistency.
Token Cleanup: In order to collect the corrected dataset, we need to identify tokens representing the end of each sentence, such as periods, question marks, and exclamation marks. In the ScienceQA dataset, a newline character often follows the ‘n’ being added after it. To reduce the burden of feedback, we simply hard-code the removal of repetitive sentences before adding the generated rationale to the prompt. We remove this ‘n’ and also ignore the backslash (\) character. For overlapping sentences, we placed each rationale of a problem into a list. If a rationale sentence was already in the list, we did not include it again during this preprocessing step.
C.3 LoRA Adapter Training
MM-CoT utilizes two identically sized models for reasoning and inference tasks. In our approach, we replace the inference model with a LoRA adapter (Figure 4), which is added to the rationale model and consists of only one-tenth of the weights.
For LoRA adapter training for ScienceQA and A-OKVQA, we use a LoRA rank of , a LoRA , and a LoRA dropout rate of . The learning rate is set to for both and . The batch size is for and for on the ScienceQA dataset. For the A-OKVQA dataset, the batch size is for both and .
C.4 Time Comparison between LoRA Adapter and Inference Model
Rather than introducing a separate model for inference, we achieve comparable performance by adding the LoRA adapter to the rationale model, while simultaneously obtaining a 2x–14x speedup in inference (ARES) compared to MM-CoT, which introduces a separate inference model. We provide the time comparison in Table 7 and Table 8.
This speed gap is mainly due to the fact that the separate inference model requires rationale generation before the inference procedure. However, the LoRA adapter directly refers to the rationale model’s latent features to derive answers, eliminating the need to generate the rationale first.
Model Rationale Inference Total 1h 32m 7m 1h 39m - 8m 8m 5h 23m 12m 5h 35m - 24m 24m
Model Rationale Inference Total 6m 2m 8m - 3m 3m 16m 3m 19m - 6m 6m
Appendix D Comparison of Generated Rationales
As mentioned in Section 2.3 and Section 4.1, because RL increases the probability of sentences receiving positive rewards and reduces the probability of sentences receiving negative rewards, the trained model often exhibits specific phenomena. It tends to generate repetitive and incomplete sentences (Table 12 and Table 13). Before the RL steps, the model couldn’t produce rationales, but after RL steps, it starts generating meaningful rationale reasoning (Table 14). Furthermore, it begins to generate reasons why other options are not the answer (Table 15).
As illustrated in Table 17, we compare the solutions from the ScienceQA original dataset, the rationales generated by the baseline model (), the rationales from the baseline model with correction feedback applied, and the rationales generated by our model (). The first example, "Which property do these three objects have in common?" illustrates that the baseline model generates incorrect rationales such as "The lemon is not (yellow)" and "All three objects are rough. The property that all three objects have in common is rough." However, when we apply correction feedback to the rationales generated by the baseline model and compare it to our proposed method, we see that our approach generates correct rationales that include the correct answer and provide explanations on why other options are not the answer. The second example also shows that our method improves rationale reasoning.
| Feedback | Prompt Structure |
| Correction Feedback |
[Prompt when Image is provided]
Your task involves reviewing a set that includes an Image, Options, Hint, Answer, and Rationales for a Question. Please follow below 7 rules. 1. Preserve any correct original rationales based on the given answer by incorporating them into the final rationale without making any alterations. 2. Preserve any original rationales that represent the starting point of thought. 3. Correct any grammatical errors or incomplete rationales based on the given information without your knowledge. 4. If there are incorrect rationales based on the given answer, please correct them without removing them based on the given information. 5. Please take into account the content of the options, hint, and answer when doing this task. 6. Fill the corrected rationales inside the {} in the final_rationale according to the given format below, without any additional explanation. 7. Return only the entire set of Rationales within curly braces ({}) below with the filled one in the step 6. Question: <Question> Options: <Choices> Hint: <Hint> Answer: <Answer> Rationales: { original_rationale:{<Rationales>} final_rationale:{} } [Prompt when no Image is provided] Your task involves reviewing a set that includes Options, Hint, Answer, and Rationales for a Question. … <same as the prompt when the image is provided> |
| A Challenging Example from A-OKVQA |
| Image |
| Question | What is the correct PSI for a bicycle tire? |
| Choices | 100-120 psi, 80-130 psi, 40-90 psi, 50-80 psi |
| A Challenging Example from A-OKVQA |
| Image |
| Question | The first two numbers of the identification tag denote what? |
| Choices | birth month, own number, birth year, birth date |
| Example of the Repetitive Sentences After the RL Stage | |
| Problem: Which tense does the sentence use? "Mona will print her name with care." | |
| Choices: [present tense, future tense, past tense] | |
| Answer: future tense | |
| Before RL | [Rationale reasoning] |
| The sentence is in future tense. You can tell because it uses will before the main verb, print. The verb tells you about something that is going to happen. | |
| After RL | [Rationale reasoning] |
| The sentence is in future tense. You can tell because it uses will before the main verb, print. The verb tells you about something that is going to happen. The verb tells you about something that is going to happen. The verb tells you about something that is going to happen. The verb ends in -s and tells you about something that is going to happen… | |
| Example of the Incomplete Sentence After the RL Stage | |
| Problem: What is the volume of a large soup pot? | |
| Choices: [7 liters, 7 milliliters] | |
| Answer: 7 liters | |
| Before RL | [Rationale reasoning] |
| The better estimate for the volume of a large soup pot is 7 liters. 7 milliliters is too little. | |
| After RL | [Rationale reasoning] |
| The better estimate for the volume of a large soup pot is- | |
| Changed Empty Rationale Reasoning Through RL | |
| Problem: Select the solid. | |
| Choices: [rain, water in a fishbowl, hammer] | |
| Answer: hammer | |
| Before RL | [Rationale reasoning] |
| <empty> | |
| After RL | [Rationale reasoning] |
| Rain is a liquid. A liquid takes the shape of any container it is in. If you put rainwater into a bucket, the rainwater will take the shape of the bucket. But the rainwater will still take up the same amount of space. The water in a fishbowl is a liquid. A liquid takes the shape of any container it is in. If you pour water from a fishbowl into a different container, the water will take the shape of that container. But the water will still take up the same amount of space. A hammer is a solid. A solid has a size and shape of its own. Many hammers are made of iron and wood.The water in a fishbowl is a liquid. A liquid takes the shape of any container it is in. If you put water from a fishbowl into a different container, the water will take the shape of that container. But the water will still take up the same amount of space. A rain is a liquid. A liquid takes the shape of any container it is in. If you put rainwater into a bucket, the rainwater will take the shape of the bucket. But the rainwater will still take up the same amount of space. A hammer is a solid. A solid has a size and shape of its own. Many hammers are made of iron and wood. Both iron and wood are solids. | |
| Example of the Process of Elimination After the RL Stage | |
| Problem: Which is a compound sentence? | |
| Choices: ["Last night we heard an owl outside in the oak tree.", | |
| "I ate all of my dinner, so I can have some pie."] | |
| Answer: "I ate all of my dinner, so I can have some pie." | |
| Before RL | [Rationale reasoning] |
| The first sentence is the compound sentence. It is made up of two simple sentences joined by a comma and the conjunction so. I ate all of my dinner, so I can have some pie. | |
| After RL | [Rationale reasoning] |
| The first sentence is the compound sentence. It is made up of two simple sentences joined by a comma and the conjunction so. I ate all of my dinner, so I can have some pie. The second sentence does not have a comma, so it is not a compound sentence. Last night we heard an owl outside in the oak tree. | |
| Comparison of Corrected Rationales After RL | |
| Image |
| Question | Will these magnets attract or repel each other? |
| Choices | repel, attract |
| After RL | Will these magnets attract or repel? To find out, look at which poles are closest to each other. The north pole of one magnet is closest to the south pole of the other magnet. Poles that are the same repel. So, these magnets will repel each other. The picture shows two magnets that are the same shape and the same size. Both magnets are the same size. So, these magnets will repel each other. |
| Corrected Rationale | The magnets will attract each other because the north pole of one magnet is closest to the south pole of the other magnet. Opposite poles of magnets attract each other in this diagram. Therefore, the correct answer is that the magnets will attract. |
| Question | How long is a pen? |
| Choices | 20 kilometers, 20 centimeters |
| After RL | The better estimate for the length of a pen is 20 centimeters. 20 kilometers is too long. 20 centimeters is too short. 20 kilometers is too long. |
| Corrected Rationale | The better estimate for the length of a pen is 20 centimeters. 20 kilometers is an unreasonable estimate, as it is far too long for a typical pen. 20 centimeters is a more accurate and reasonable estimate for the length of a standard pen. |
| Image |
| Question | What is the capital of South Carolina? |
| Choices | Columbia, Montgomery, Charleston, Harrisburg |
| After RL | Columbia is the capital of South Carolina.nThe city of Columbia is Charleston, South Carolina. Charleston is the smallest city in South Carolina.nHayeburg is the smallest city in South Carolina. Montgomery is the smallest city in South Carolina. But the city of Columbia is home to many cultural and historical landmarks. |
| Corrected Rationale | Columbia is the correct answer as it is the capital city of South Carolina. It is the largest city in the state and serves as the seat of government. The other options, Montgomery (the capital of Alabama), Charleston (a city in South Carolina), and Harrisburg (the capital of Pennsylvania), are not the capital of South Carolina. |
| Comparison of Generated Rationales | |
| Image |
| Question | Which property do these three objects have in common? |
| Choices | rough, yellow, sticky |
Original
Solution |
Look at each object. For each object, decide if it has that property. A rough object feels scratchy when you touch it. The rain boots and the rubber duck are not rough. A sticky object can attach or stick to other things. The rain boots and the rubber duck are not sticky. Yellow is a color. This color is yellow. All three objects are yellow. The property that all three objects have in common is yellow. |
| Look at each object. For each object, decide if it has that property. A sticky object can attach or stick to other things. The rubber duck is not sticky. Yellow is a color. This color is yellow. The rubber duck and the rain boots are yellow, but the lemon is not. A rough object feels scratchy when you touch it. All three objects are rough. The property that all three objects have in common is rough. | |
Correction
feedback on |
The property that all three objects have in common is yellow. The rubber duck and the rain boots are both yellow, and the lemon is also yellow. Although the objects may have other properties like roughness, the common property among them is the color yellow. |
| (Ours) | The three objects shown - rubber duck, rain boots, and lemon - all share the common property of being yellow in color. The rubber duck is made of a smooth, non-sticky material, the rain boots are also a yellow-colored object, and the lemon is also a yellow-colored liquid. Therefore, the correct answer is that the common property among these three objects is that they are yellow. The other options, rough and sticky, do not apply to all three objects, as the rubber duck is not rough, and the rain boots are not sticky. |
| Question | Which word would you find on a dictionary page with the following guide words? lever - loan |
| Choices | lamb, liquid |
Original
Solution |
Put the words in alphabetical order. Since liquid is between the guide words lever - loan, it would be found on that page. |
| Put the words in alphabetical order. Since lamb is between the guide words lever - loan, it would be found on that page. | |
Correction
feedback on |
The word "liquid" would be found on the dictionary page with the guide words "lever - loan" because it falls alphabetically between those two words. |
| (Ours) | The word "liquid" would be found on the dictionary page with the guide words "lever - loan" because it falls alphabetically between those two words. Therefore, the word "liquid" would be the word found on that dictionary page. |